solve using the elimination method.
3x+5y=4
-2x+3y=10

Answer by london maths tutor(243)

3x+5y=4 ... equation 1
-2x+3y=10 ... equation 2
equation 1 x 2 :
3x(2) + 5y(2) = 4(2)
6x + 10y = 8
equation 2 x 3 :
-2x(3)+3y(3) = 10(3)
-6x + 9y = 30
Equation 1 + Equation 2
6x + 10 y + [-6x + 9y] = 8 + 30
19 y = 38
y = 38/19
y = 2
Put y = 2 into 3x + 5y = 4 to solve x
3x + 5(2) = 4
3x + 10 =4
3x = - 6
x = - 6/3
x = - 2
Answer: x = -2 and y = 2